OpenStudy (anonymous):

write the slope-intercept equation for the line that passes through (-7,6) and is perpendicular to -7x+9y=-2 please show all work

OpenStudy (zepp):

Use this guide right there: http://openstudy.com/users/zepp#/updates/4fa3474fe4b029e9dc34125e And tell me if you still don't understand.

OpenStudy (anonymous):

im still lost

OpenStudy (zepp):

Okay, let's transform that into the general form: -7x+9y=-2 -7x+9y+2=0 From my tutorial, the slope of this function is \(\frac{-A}{B}=\frac{-(-7)}{9}=\frac{7}{9}\) Now we are looking for the slope of a function that's perpendicular to this one; so we take this slope, and twist it, change its sign (In fact, reciprocal of the slope with a different sign): \(-\frac{9}{7}\)

OpenStudy (zepp):

Now we have a slope \(-\frac{9}{7}\) and a point (-7,6) Let's take the slope-intercept equation: \(y=mx+b\) Plug in the point (-7,6) and the slope \(6=(-\frac{9}{7})(-7)+b\) and find b. \(6=9+b\) \(-3=b\) So the equation would be \(y=-\frac{9}{7}x-3\)

OpenStudy (anonymous):

thank you, you were very helpful. I will keep some notes to help me out

OpenStudy (zepp):

You are very welcome! :)